Find a parametrization of cylinder. x^2+〖(y-4)〗^2=16, 0≤z≤5
1
Expert's answer
2021-07-13T05:05:22-0400
"x=4\cos \theta, y-4=4\sin \theta"

The cylinder "x^2+(y-4)^2=16, 0\leq z\leq5" is parametrized as


"\bold{r}(\theta, z)=\langle4\cos\theta, 4+4\sin\theta,z\rangle,"




"\theta\in(0, 2\pi], z\in[0,5]"